OpenStudy (anonymous):

Solve:

OpenStudy (sasogeek):

still typing... i wonder how long the question is :P wasiqss, get ready

OpenStudy (wasiqss):

lol saso i actually posted n L opital rule ques can u solve

OpenStudy (anonymous):

\[\left(\begin{matrix}x^2 \over x-1\end{matrix}\right) \div\]\[\left(\begin{matrix}x^2 \over y-xy\end{matrix}\right)\]

OpenStudy (anonymous):

For some reason i had a really hard time typing that =/

OpenStudy (wasiqss):

y is anwer

OpenStudy (sasogeek):

solve this? i think u mean simplify... right? otherwise solve for ... ?

OpenStudy (anonymous):

lol yeah simplify

OpenStudy (wasiqss):